# A Hierarchical Bin Packing Framework with Dual Manipulators via Heuristic Search and Deep Reinforcement Learning

### v1

@@ Abstract
Index Terms:
Bin packing, dual manipulators, warehouse automation, hierarchical planning
@@ I Introduction
BPP includes offline, online, and semi-online variants, depending on the availability of item information [11]. While the offline BPP assumes full knowledge of all items, real-world applications typically involve online or semi-online settings, where the information of items are revealed sequentially. Semi-online BPPs are more common in industrial contexts such as logistics and manufacturing. In such settings, strategies for managing uncertainty, such as repacking, can improve adaptability to changes in item arrivals. BPP is also categorized by packing dimension (e.g., 2D, 3D) [12]. Although there has been a line of research on both 2D BPP and 3D BPP, 2D BPP remains underexplored, particularly in achieving near-optimal packing through rotation and/or repacking.
Most existing BPP methods assume a single-robot setting, although coordinating multiple manipulators can significantly reduce execution time through parallelization. Coordinating multiple robots for BPP remains computationally challenging due to the inherent complexity of multi-robot task allocation and multi-arm motion planning [13, 14, 15, 16].
@@
  • A heuristic search algorithm that supports a wide range of 2D BPP scenarios, including offline, semi-online, and online settings, with varying levels of item accessibility and information, and even multi-manipulator configurations.
@@ II Related Works
Prior work has addressed offline and online 2D BPPs, with recent advances extending to semi-online settings and incorporating rotation and repacking techniques. This section reviews related research on these BPP variants and strategies for rotation and repacking. We also discuss approaches for coordinating multi-robot operations in shared workspaces.
@@
To fill this gap, we propose a framework for 2D BPP under realistic multi-robot settings. Our method considers item accessibility through forward simulation to ensure feasibility, and integrates coordinated task planning and repacking within a unified hierarchical structure. Leveraging a dual-manipulator system, we enable parallel execution while ensuring packing decisions are both feasible and optimized for bin utilization.
@@ III Problem Formulation
(a) The workspace and cache of the dual manipulators
@@
Figure 2: Illustration of the Bin Packing system. (a) The manipulator is denoted by $m$, $\mathcal{Z}$ represents the workspace of the manipulator, and $C$ is the temporary storage area for unpacked items. (b) The bin is an open cuboid, and the next item $o$ is a solid cuboid. (c) The bin is modeled as a padded binary grid matrix $B$, where each cell corresponds to a low-level position action indexed by its coordinate in the image coordinate system. The values increase row-wise from left to right and top to bottom, with the origin located at the top-left corner. The item is encoded as the rotated vector $\mathbf{l}(\phi)$, and a checkerboard-patterned pixel indicates the top-left corner of the placed item.
@@
As shown in Figs. 2(a) and 2(b), we consider a 2D BPP environment where the bin is modeled as an open cuboid and items as closed cuboids, all projected onto the $xy$-plane. We assume up to two robotic manipulators, denoted by $m_{i}$ for $i\in\{1,2\}$, each with a reachable workspace $\mathcal{Z}_{i}\subset\mathcal{W}\subseteq\mathbb{R}^{3}$. Items arrive sequentially via a conveyor belt (or other similar facilities), which advances as items are removed. The bin and each item $o$ are discretized into grid cells: the bin is a $W\times H$ grid, and item $o$ occupies $w_{o}\times h_{o}$ cells, where $W,H,w_{o},h_{o}\in\mathbb{Z}^{+}$. The bin occupancy is represented by a padded binary matrix $B\in\{0,1\}^{(W+2)\times(H+2)}$, where 00 and $1$ indicate free and occupied cells. The outermost rows and columns are padded with $1$s to represent the bin boundaries (see Fig. 2(c)). The rotated size vector $\mathbf{l}_{o}(\phi)=[l_{x}^{(o,\phi)},l_{y}^{(o,\phi)}]^{T}$ defines the dimensions of item $o$ after rotation by $\phi\in\{0,\frac{\pi}{2}\}$. For example, if $w_{o}=3$ and $h_{o}=1$, then $\mathbf{l}_{o}(0)=[3,1]^{T}$ and $\mathbf{l}_{o}(\frac{\pi}{2})=[1,3]^{T}$. We denote $\mathcal{N}$ as items on the conveyor, $\mathcal{C}$ as unpacked items, and $\mathcal{I}$ as packed items. Each manipulator $m_{i}$ has a buffer space $C_{i}\subset\mathcal{Z}_{i}$ to temporarily store items in $\mathcal{C}$.
The High-level MDP
@@
where $B$ is the current bin occupancy and $\mathbf{l}_{o}(\phi)$ is the rotated dimension of $o$. Given $s_{\text{low}}^{(o,\phi)}$, the low-level action $a_{\text{low}}^{(o,\phi)}\in\mathcal{A}_{\text{low}}$ specifies the position for placing the top-left corner of $o$ in the bin. The high-level MDP evaluates both orientations and selects one along with its corresponding position to complete the pack primitive. The bin position is represented by an index from 00 (top-left) to $W\cdot H-1$ (buttom-right), following a row-wise order as illustrated in Fig. 2(c). Given a position $(x_{o},y_{o})$ for item $o$, the low-level action is defined as:
@@ IV-B Interaction Between High-level and Low-level
This leads to a new vertex $v^{\prime}=\{B^{\prime},\mathcal{N}^{\prime},\mathcal{I}^{\prime},\mathcal{C}^{\prime}\}\in\mathcal{V}$ in the tree, where $B^{\prime}$ is updated via (1), and item $o$ is moved from $\mathcal{N}$ or $\mathcal{C}$ to $\mathcal{I}$ [^1] Item $o$ is removed from $\mathcal{N}$ or $\mathcal{C}$, whichever contains it, and added to $\mathcal{I}$. If $a_{\text{low}}^{(o,\phi)}$ is the no-position action (2), no child is generated and $v$ becomes a leaf vertex. A leaf represents a state where either (i) no items remain in $\mathcal{N}\cup\mathcal{C}$, or (ii) $o$ cannot be placed due to collisions with packed items in $\mathcal{I}$.
@@ IV-C1 High-Level Search Algorithm
1:  $v_{\text{root}}\leftarrow s_{\text{high}}$ //Initialization
@@
[…]
@@ IV-C2 Tree Construction
1:  $\mathcal{O}\leftarrow();$ $n=0?\ stop=\text{\bf{true}}:stop=\text{\bf{false}}$ //Initialization
@@
[…]
@@ IV-C3 Action Generation and Forward Simulation
Once the tree is fully expanded, actions are generated and forward simulation is performed for each $\tilde{\chi}\in\mathbb{X}$, constructed by Alg. 2. Through forward simulation, the score $\mu(\tilde{\chi})$ is calculated from an initial value of zero. Each element of $\tilde{\chi}$ is a tuple $(o,\phi,a^{(o,\phi)}_{\text{low}},d(v^{\prime}))$. Tuples with valid positional actions are sequentially converted into pack$(o,\phi,x_{o},y_{o})$ commands. If $o$ is an unpacked or accessible item, the pack action is executed in the simulation, and the corresponding low-level reward from $a^{(o,\phi)}_{\text{low}}$ is added to $\mu(\tilde{\chi})$. In conveyor-based environments, if an item is removed from $\mathcal{N}$ via pack, subsequent items advance, potentially making previously inaccessible items accessible. If the action is a no-position action or the item remains inaccessible, the item is skipped in the simulation. The simulation terminates when either all accessible items result in no-position actions or all tuples in $\tilde{\chi}$ have been processed. Then, $util(\tilde{\chi})$, the final utilization ratio of $\tilde{\chi}$, is computed.
As the forward simulation progresses, pack actions are appended to the corresponding high-level action $a_{\tilde{\chi}}$. All pack actions for unpacked items are included. For accessible items not yet packed, at most as many pack primitives are added as the number of available robots. Other accessible items with valid positional actions are executed in the simulation but excluded from $a_{\tilde{\chi}}$ at the current step. Among accessible items, inclusion in $a_{\tilde{\chi}}$ is based on the depth value $d(v^{\prime})$. In a dual-arm setting, both robots must pick items simultaneously. The item with the smallest $d(v^{\prime})$ accessible by at least one robot is selected first. If only one robot can access it, the other is assigned the item with the next smallest $d(v^{\prime})$ it can access. If both robots can access the same item, the second smallest $d(v^{\prime})$ item accessible by either is selected next. If the next conveyor step yields no packable items, a terminate primitive is appended to $a_{\tilde{\chi}}$.
After the forward simulation, the sequence with the highest score $\mu(\tilde{\chi})$ is selected. If multiple sequences yield the same score, the one with the highest utilization ratio $util(\tilde{\chi})$ is chosen. If a tie still remains, the sequence with the smallest total $d(v^{\prime})$ for items included in the high-level action is selected, as prioritizing items with smaller depth values improves adaptability to unknown future arrivals, as illustrated in Fig. 6.
@@ IV-C4 Repacking Solution Search
[…]
@@ IV-D Task Allocation and Atomic Action Sequencing
TABLE I: Scenario configurations with varying robot and item setups.
@@ V Experiments
We evaluate our hierarchical bin packing framework using a hybrid architecture, integrating task planning, DRL, and motion planning modules across two systems. The algorithm runs on a machine with an AMD Ryzen 7 9700X CPU, 64 GB RAM, and a GeForce RTX 4080 SUPER GPU. Simulation and motion planning are executed on a separate system with an AMD Ryzen 7 5800X CPU, 32 GB RAM, and a GeForce RTX 3070 Ti GPU. Modules communicate via ROS1 [26] and TCP/IP to synchronize the interaction between the planner and simulation. The DRL agent is implemented in PyTorch [27], and the simulation environment is developed in Unity [28], providing physical dynamics. All training and evaluation experiments are conducted in a $10\times 10$ bin environment.
@@ V-A Dataset
In the Random instance, items are generated by sampling the width and height of each item uniformly, with each dimension limited to at most half the bin size. A large number of sets are included to ensure sufficient diversity and test coverage.
The 100% set instances are constructed to train the RL agent and evaluate the algorithm. It consists of item sets where full bin utilization (100%) is achievable. These sets enable the agent to learn optimal packing strategies and serve as a benchmark to test whether repacking can fully utilize the bin space. To construct each set, item dimensions are randomly sampled and placed at feasible locations to ensure compact, valid placements. Oversized items are avoided by considering remaining space during sampling. Once the bin is fully packed, the item order is shuffled to simulate online arrival. In total, 9.2 million item sets are generated across 100% set instances: 7.2 million (78%) are used to train the A3C worker networks, and 2 million (22%) are reserved exclusively for evaluating the global network.

### v2

@@ I Introduction
BPP includes offline, online, and semi-online variants, depending on the availability of item information [11]. While the offline BPP assumes full knowledge of all items, real-world applications typically involve online or semi-online settings, where the information of items are revealed sequentially. Semi-online BPPs are more common in industrial contexts such as logistics and manufacturing. In typical warehouse or inspection setups, a camera or barcode reader is installed above a conveyor belt, allowing partial observation of incoming items within a fixed field of view. In such cases, the system can access limited item information (e.g., size, orientation, or label data) before packing, which aligns with the assumptions of the 2D online and semi-online settings. In such settings, strategies for managing uncertainty, such as repacking, can improve adaptability to changes in item arrivals. BPP is also categorized by packing dimension (e.g., 2D, 3D) [12]. Although there has been a line of research on both 2D BPP and 3D BPP, many real-world applications are effectively 2D due to physical or operational constraints. For example, fragile or breakable items such as glass panels and ceramics cannot be stacked; inspection or vision-based operations often require items to remain within a single camera view; and layout optimization on flat materials such as textiles, leather, or printed circuit boards (PCBs) is inherently 2D. In addition, certain industrial containers, such as interlocking plastic crates (e.g., milk crates), and uniformly shaped items that can be stacked in shallow multi-layer configurations, also fit naturally within a 2D formulation. In these environments, the packing task is dominated by footprint efficiency rather than vertical stacking. Motivated by these practical use cases, this work focuses on the 2D online and semi-online bin packing problem.
Most existing BPP methods assume a single-robot setting, although coordinating multiple manipulators can significantly reduce execution time through parallelization. Coordinating multiple robots for BPP remains computationally challenging due to the inherent complexity of multi-robot task allocation and multi-arm motion planning [13, 14, 15, 16]. While asynchronous or semi-synchronous operation can potentially improve execution efficiency, these approaches considerably increase the complexity and computation time of task and motion planning required to ensure safe and collision-free coordination. Therefore, the proposed framework adopts a synchronous operation, which provides predictable coordination between manipulators, simplifies collision avoidance, and reduces planning complexity, which, makes it more practical for real-world deployment.
@@
  • A heuristic search algorithm that integrates robot accessibility and multi-manipulator coordination into the policy generation process, enabling feasible and optimized packing decisions under offline, semi-online, and online conditions with varying perception ranges.
@@ II Related Works
Prior work has addressed offline and online 2D BPPs, with recent advances extending to semi-online settings and incorporating rotation and repacking techniques. This section reviews related research on these BPP variants and strategies for rotation and repacking. We also discuss approaches for coordinating multi-robot operations in shared workspaces. Table I summarizes representative 2D and 3D BPP studies, focusing on their problem settings and robotic implementation aspects such as repacking, multi-bin handling, and dual-arm coordination.
TABLE I: Comparison of representative studies in 2D and 3D bin packing settings.
Characteristic 2D 3D
Offline / Semi-online Ours, [3] [6, 18, 8, 19, 20, 21, 22, 23, 24, 25]
Online Ours,[7, 26, 4, 5] [19, 20, 21, 22, 23]
Repacking Ours [22, 23, 24]
Multi-bin [6, 18, 19, 24]
Robotic Implementation Ours [18, 19, 20, 22, 23, 24, 25]
Dual-arm Allocation per bin Ours
@@
Recent studies have also explored multi-arm in bin packing scenarios.  [24] proposes a deep reinforcement learning framework for online and concurrent 3D bin packing with bin replacement, in which multiple bins are processed in parallel to improve throughput. Their work focuses on optimizing packing efficiency across multiple containers, rather than coordinating multiple manipulators within a single bin. On the other hand,  [25] develops a dual-arm robotic packing framework emphasizing joint motion coordination and trajectory planning for cooperative manipulation of a single object. While both studies address related challenges in robotic packing, their problem formulations differ from ours: we focus on dual-manipulator cooperation for packing multiple objects within a single bin under online and semi-online settings, where task allocation and placement decisions are jointly optimized through a hierarchical algorithm.
@@
To fill this gap, we propose a framework for 2D BPP under realistic dual-arm settings. Unlike previous studies that determine object placement and orientation before considering robotic execution, our method jointly plans the packing policy while accounting for the reachability of both single and dual manipulators. This integrated formulation ensures that packing decisions are physically feasible for robot execution while maintaining near-optimal packing efficiency. Moreover, the framework synergistically integrates a low-level RL policy with a high-level tree search, combining the flexibility of learning-based decision making with the reliability of search-based reasoning.
@@ III Problem Formulation
(a) The task space and cache of the dual manipulators
@@
Figure 2: Illustration of the Bin Packing system. (a) The manipulator is denoted by $m$, $\mathcal{Z}$ represents the task space of the manipulator, and $C$ is the temporary storage area for unpacked items. (b) The bin is an open cuboid, and the next item $o$ is a solid cuboid. (c) The bin is modeled as a padded binary grid map $B$, where each cell corresponds to a low-level position action indexed by its coordinate in the image coordinate system. The values increase row-wise from left to right and top to bottom, with the origin located at the top-left corner. The item is encoded as the rotated vector $\mathbf{l}(\phi)$, and a checkerboard-patterned pixel indicates the top-left corner of the placed item.
@@
As shown in Figs. 2(a) and 2(b), we consider a 2D BPP environment where the bin is modeled as an open cuboid and items as closed cuboids, all projected onto the $xy$-plane. We assume up to two robotic manipulators, denoted by $m_{i}$ for $i\in\{1,2\}$, each with a task space $\mathcal{Z}_{i}\subset\mathcal{W}\subseteq\mathbb{R}^{3}$. Items arrive sequentially via a conveyor belt (or other similar facilities), which advances as items are removed. The bin and each item $o$ are discretized into grid cells: the bin is a $W\times H$ grid, and item $o$ occupies $w_{o}\times h_{o}$ cells, where $W,H,w_{o},h_{o}\in\mathbb{Z}^{+}$. The bin occupancy is represented by a padded binary map $B\in\{0,1\}^{(W+2)\times(H+2)}$, where 0 and $1$ indicate free and occupied cells. The outermost rows and columns are padded with $1$s to represent the bin boundaries (see Fig. 2(c)). The rotated size vector $\mathbf{l}_{o}(\phi)=[l_{x}^{(o,\phi)},l_{y}^{(o,\phi)}]^{T}$ defines the dimensions of item $o$ after rotation by $\phi\in\{0,\frac{\pi}{2}\}$. For example, if $w_{o}=3$ and $h_{o}=1$, then $\mathbf{l}_{o}(0)=[3,1]^{T}$ and $\mathbf{l}_{o}(\frac{\pi}{2})=[1,3]^{T}$. We denote $\mathcal{N}$ as items on the conveyor, $\mathcal{C}$ as unpacked items, and $\mathcal{I}$ as packed items. Each manipulator $m_{i}$ has a buffer space $C_{i}\subset\mathcal{Z}_{i}$ to temporarily store items in $\mathcal{C}$.
III-A The High-level MDP
@@ III-B The Low-level MDP
where $B$ is the current bin occupancy and $\mathbf{l}_{o}(\phi)$ is the rotated dimension of $o$. Given $s_{\text{low}}^{(o,\phi)}$, the low-level action $a_{\text{low}}^{(o,\phi)}\in\mathcal{A}_{\text{low}}$ specifies the position for placing the top-left corner of $o$ in the bin. The high-level MDP evaluates both orientations and selects one along with its corresponding position to complete the pack primitive. The bin position is represented by an index from 0 (top-left) to $W\cdot H-1$ (buttom-right), following a row-wise order as illustrated in Fig. 2(c). Given a position $(x_{o},y_{o})$ for item $o$, the low-level action is defined as:
@@ IV-B Interaction Between High-level and Low-level
This leads to a new vertex $v^{\prime}=\{B^{\prime},\mathcal{N}^{\prime},\mathcal{I}^{\prime},\mathcal{C}^{\prime}\}\in\mathcal{V}$ in the tree, where $B^{\prime}$ is updated via (1), and item $o$ is moved from $\mathcal{N}$ or $\mathcal{C}$ to $\mathcal{I}$. [^1] Item $o$ is removed from $\mathcal{N}$ or $\mathcal{C}$, whichever contains it, and added to $\mathcal{I}$. If $a_{\text{low}}^{(o,\phi)}$ is the no-position action (2), no child is generated and $v$ becomes a leaf vertex. A leaf represents a state where either (i) no items remain in $\mathcal{N}\cup\mathcal{C}$, or (ii) $o$ cannot be placed due to collisions with packed items in $\mathcal{I}$.
@@ IV-C1 High-Level Search Algorithm
1:  $v_{\text{root}}\leftarrow s_{\text{high}}$ .//Initialization
@@
If no valid solution is found ($\mathbb{X}=\emptyset$), Alg. 1 sets $a_{\text{high}}$ to terminate to end the episode, and assigns utilutil, the highest achievable utilization ratio identified so far, as the current value (lines 35). Otherwise, each sequence $\tilde{\chi}\in\mathbb{X}$ is evaluated, and the candidate high-level action set $A$ is generated through SimulationAndGeneration (line 7). Then, BestActionSelection selects the best sequence $\chi^{*}$ and sets its action as $a_{\text{high}}$ (line 8). Both procedures are detailed in Sec. IV-C3. After $\chi^{*}$ is determined, utilutil is assigned as $\mathit{util}(\chi^{*})$, the utilization ratio computed from $\chi^{*}$ (line 9).
@@ IV-C2 Tree Construction
1:  $\mathcal{O}\leftarrow();$ $n=0?\ stop=\text{\bf{true}}:stop=\text{\bf{false}}$ .//Initialization
@@
10:   if requireFullPack and $(a_{\text{low}}^{(o,0)},a_{\text{low}}^{(o,\pi/2)}\text{ are NoPositionAction})$ then
@@ IV-C3 Action Generation and Forward Simulation
[…]
As the forward simulation progresses, pack actions are appended to the corresponding high-level action $a_{\tilde{\chi}}$. All pack actions for unpacked items are included. For accessible items not yet packed, at most as many pack primitives are added as the number of available robots. Other accessible items with valid positional actions are executed in the simulation but excluded from $a_{\tilde{\chi}}$ at the current step. Among accessible items, inclusion in $a_{\tilde{\chi}}$ is based on the depth value $d(v^{\prime})$. In the proposed synchronized dual-arm setting, two objects are selected and assigned to the respective robots. The item with the smallest $d(v^{\prime})$ accessible by at least one robot is selected first. If only one robot can access it, the other is assigned the item with the next smallest $d(v^{\prime})$ it can access. If both robots can access the same item, the second smallest $d(v^{\prime})$ item accessible by either is selected next. If the next conveyor step yields no packable items, a terminate primitive is appended to $a_{\tilde{\chi}}$.
[…]
@@ IV-C4 Repacking Solution Search
1:  $a^{*}\leftarrow(),a\leftarrow(),RepackSuccess$ $\leftarrow\textbf{false}{}$ .//Initialization
@@ IV-D Task Allocation and Atomic Action Sequencing
TABLE II: Scenario configurations with varying robot and item setups.
@@ V Experiments
We evaluate our hierarchical bin packing framework using a hybrid architecture, integrating task planning, DRL, and motion planning modules across two systems. The algorithm runs on a machine with an AMD Ryzen 7 9700X CPU, 64 GB RAM, and a GeForce RTX 4080 SUPER GPU. Simulation and motion planning are executed on a separate system with an AMD Ryzen 7 5800X CPU, 32 GB RAM, and a GeForce RTX 3070 Ti GPU. Modules communicate via ROS1 [28] and TCP/IP to synchronize the interaction between the planner and simulation. The DRL agent is implemented in PyTorch [29], and the simulation environment is developed in Unity [30], providing physical dynamics. All training and evaluation experiments are conducted in a $10\times 10$ bin environment. Considering the placement margin, robot positioning accuracy, and camera perception error, the 10 × 10 grid resolution is sufficiently precise for stable packing and is therefore adopted as the representative configuration for experimental validation.
@@ V-A Dataset
In the Random instance, items are generated by independently sampling each dimension uniformly: $w_{o}\sim\mathcal{U}(1,\lfloor W/2\rfloor)$ and $h_{o}\sim\mathcal{U}(1,\lfloor H/2\rfloor)$, ensuring no single item exceeds half the bin size in either dimension.
100% Set instances consist of item sets where full bin utilization (100%) is achievable, enabling the agent to learn optimal packing strategies and serving as a benchmark to test whether repacking can fully utilize the bin space. The construction procedure is detailed in Alg. 4. Item dimensions are sampled from Gaussian-weighted probabilities (standard deviation $\sigma=2$) over width and height, which reduces the probability of generating items with dimension 1 in early stages. This design prevents an overabundance of such small or elongated items that would inevitably arise later when filling remaining gaps. For each sampled item, we compute edge-contact rewards for all feasible placement positions and greedily select the position that maximizes this reward, with random tie-breaking when multiple positions share the maximum value. Oversized items that cannot fit in the remaining space are rejected and resampled. Once the bin is full or only one cell remains empty, the process is terminated. In the case where one cell remains empty, a $1\times 1$ item is added to achieve exact 100% fill. Finally, the item order is shuffled to simulate online arrival. In total, 9.2 million item sets are generated: 7.2 million (78%) are used to train the A3C worker networks, and 2 million (22%) are reserved exclusively for evaluating the global network.
Algorithm 4 100% Set Instance Generation
0:  Bin size $W\times H$, Gaussian std $\sigma=2$
0:  Item set $\mathcal{S}$ with 100% bin utilization
1:  $\mathcal{S}\leftarrow\emptyset$, $\text{area}\leftarrow 0$
2:  Initialize empty bin $B$ of size $W\times H$
3:  $\max_{w}\leftarrow 0$, $\max_{h}\leftarrow 0$
4:  $p_{w}\leftarrow\text{GaussianProb}(W,\sigma)$
5:  $p_{h}\leftarrow\text{GaussianProb}(H,\sigma)$
6:  while $\text{area}<W\times H-1$ do
7:   repeat
8:    repeat
9:     Sample $w_{o}\sim p_{w}$, $h_{o}\sim p_{h}$
10:    until $(w_{o}+h_{o}\leq W+H-\max_{w}-\max_{h})$ AND    $(\text{area}+w_{o}\cdot h_{o}\leq W\times H)$
11:    Initialize $R[y,x]\leftarrow$-$\infty$ for all $(x,y)$
12:    Compute $R[y,x]\leftarrow$ $\text{EdgeContact}(B,x,y,w_{o},h_{o})$ for feasible $(x,y)$
13:   until $\max_{(x,y)}R[y,x]\geq 0$
14:   $(x^{*},y^{*})\leftarrow\arg\max_{(x,y)}R[y,x]$ with random tie-breaking
15:   Place item $(w_{o},h_{o})$ at position $(x^{*},y^{*})$ in bin $B$
16:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{(w_{o},h_{o})\}$
17:   $\text{area}\leftarrow\text{area}+w_{o}\cdot h_{o}$
18:   $\max_{w}\leftarrow\max(\max_{w},w_{o})$, $\max_{h}\leftarrow\max(\max_{h},h_{o})$
19:  end while
20:  if $\text{area}=W\times H-1$ then
21:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{(1,1)\}$
22:  end if
23:  Randomly shuffle $\mathcal{S}$
24:  return $\mathcal{S}$